OpenStudy (anonymous):

Write an equation of a line parallel to y=1/4x - 6 in slope-intercept form that passes through the point (-1,5)

OpenStudy (anonymous):

y= -4x + 5 y= 1/4x - 1 y= -4x - 1.25 y= 1/4x + 5.25

OpenStudy (anonymous):

help?!

OpenStudy (myth1111):

Well, if the equations are parallel, the slope should be the same, so it is either B or D.

OpenStudy (anonymous):

okay, which one do you think it is,.. I think maybe it could be D?!

OpenStudy (myth1111):

You are correct.

OpenStudy (anonymous):

thanks:)